# Leptogenesis from magnetic helicity of gauged $\text{U}(1)_{B-L}$

### v1

@@ IV Model with Dirac neutrinos
As we have discussed in the previous section, we need two requirements to generate the baryon asymmetry of the universe. The first requirement is the washout of the $B-L$ charges in the standard model sector and the chiral sector. The second requirement is the decoupling of the interactions before the decay of the magnetic helicity. As the Yukawa interaction is decoupled, the first requirement is not satisfied, because the Lagrangian respects the global $B-L$ symmetry. However, the Yukawa interaction is not the only possible interaction between the right-handed neutrinos and the standard model leptons. For example, we can introduce higher-dimensional operators between the right-handed neutrinos and the standard model leptons. The most relevant operator consistent with the symmetries is a dimension-6 operator,
[…]
@@
where $\Lambda$ is the cutoff scale. In general, the cutoff scale is expected to be the Planck scale, $\Lambda\sim M_{p}$. If we introduce an additional heavy scalar field, $H^{\prime}$, with the same quantum numbers as the Higgs field, for example, the cutoff scale can be much smaller than the Planck scale. Another possibility, as seen in some models called the Dirac seesaw models (e.g., Ref. Roncadelli and Wyler (1983); Ma and Srivastava (2015)), involves the introduction of new heavy particles that equilibrate the $B-L$ charges in the standard model sector and chiral sector. As these models introduce additional charged particles and the dynamics of the $B-L$ gauge field change, for simplicity, we focus on the simplest model with the dimension-6 operator. With the dimension-6 operator, the $B-L$ charges in the standard model sector and the chiral sector are at least partially equilibrated. Let $c$ denote the fraction of the $B-L$ charge washed out; $c=1$ if the $B-L$ charges are totally equilibrated.
@@
Figure 4: The excluded region of the parameter space in terms of the $B-L$ asymmetry and the Hubble constant at the end of the inflation. The left panel is for $g_{B-L}=0.1$ and the right panel is for $g_{B-L}=0.01$. The blue region is the excluded region by the energy density of the magnetic field and the generated Fermions. The green region is the excluded region by the requirement of the magnetic Reynolds number, $R_{m}<1$. We also indicate the contour with $R_{e}=1$ by the green dotted line. The region left to the green dotted line is the region where the kinetic Reynolds number is smaller than unity. The purple, brown, light blue, and yellow dashed lines are the contours of $\xi=5,10,15,20$, respectively.
@@
In Fig. 4, we show the excluded region of the parameter space in terms of the $B-L$ asymmetry and the Hubble constant at the end of the inflation, assuming that the magnetic helicity is generated during the axion inflation as discussed in Sec. II. The blue region is the excluded region where the energy density of the magnetic field and the generated asymmetric fermion are larger than $1\,\%$ of the total energy density of the universe at the end of the inflation. The green region is the excluded region by the requirement of the magnetic Reynolds number, $R_{m}\lesssim 1$. We also show the contour with $R_{e}=1$ by the green dotted line as a reference. The region left to the green dotted line is the region where the kinetic Reynolds number is smaller than unity.
@@
From the figure, we can see that we need a larger $B-L$ helicity for the magnetic field not to dissipate for a conservative estimate of the magnetic Reynolds number adopted in Sec. II. This requires that the fraction of the $B-L$ charge washed out, $c$ is much smaller than unity, as $c\simeq\frac{79}{28}\eta_{0}/\eta_{\cal H}$ to explain the baryon asymmetry of the universe, assuming that the anomalous charge cancellation does not occur. Alternatively, we can dilute the baryon asymmetry by the entropy production after the decay of the magnetic helicity. Another possibility is to consider the case where the helicity is decaying by the dissipation or the anomalous charge cancellation, but the final $B-L$ charge is determined by the balance between the decay of the helicity and the washout process. Once more, since the condition adopted here for the magnetic field not to dissipate is a rough estimate, the possibility that the magnetic field survives at smaller $\eta_{H}$ is not completely ruled out Domcke et al. (2023). Determining the parameter space for the successful baryogenesis scenario requires more detailed analyses of the helicity evolution and decay by MHD as well as fermion production, which are beyond the scope of this paper. To conclude this section, we note that the generation of the baryon asymmetry of the universe in the model with Dirac neutrinos can be seen as an implementation of Dirac leptogenesis Dick et al. (2000); Murayama and Pierce (2002). In Dirac leptogenesis, the decay of a heavy particle produces a $B-L$ asymmetry in both the standard model sector and the right-handed neutrino sector. In our scenario, this asymmetry is generated through the decay of the magnetic helicity of the $B-L$ gauge field.
@@ V Discussion and Conclusion
In this paper, we have proposed a new scenario to generate the baryon asymmetry of the universe. We have discussed the decay of the magnetic helicity of the $B-L$ gauge field, which is generated, e.g., during axion inflation, after the $B-L$ gauge symmetry is Higgsed. The net $B-L$ asymmetry in the standard model sector is the combination of the asymmetry generated during inflation, which is subject to the washout effect, and the asymmetry generated by the $B-L$ helicity decay. For both the seesaw model with Majorana neutrinos and the model with Dirac neutrinos, we have investigated if the baryon asymmetry of the universe can be generated in the model through our mechanism. We have shown that the observed baryon asymmetry can be generated in the model with Dirac neutrinos if the effect of the washout process is decoupled before the decay of the magnetic helicity. For the model with Majorana neutrinos, we have shown that the baryon asymmetry of the universe cannot be generated in the minimal seesaw model with the second-order phase transition, although the baryon asymmetry can be generated in more complicated models.
While our discussion on the helicity evolution is based on order-of-magnitude estimates from MHD, the fundamental principles of our scenario remain robust. These principles include the conservation of $B-L$ magnetic helicity, the equilibration and subsequent decoupling of $B-L$ charges between the standard model sector and the chiral sector, and the decay of magnetic helicity following the $B-L$ phase transition. However, to precisely quantify the baryon asymmetry of the universe, detailed analytical formulation and numerical simulations of helicity decay, particularly near the phase transition, are essential.
@@
Finally, in this paper, we have ignored the effect of the $\text{U}(1)_{Y}$ gauge field and and its dynamics. For example, the kinetic mixing between the $B-L$ gauge field and the $\text{U}(1)_{Y}$ can generate the helicity of the $\text{U}(1)_{Y}$ gauge field Kamada et al. (2018). Also, if the generated $B-L$ charge is large and the temperature is low enough, the chiral plasma instability may generate the helicity of the $\text{U}(1)_{Y}$ gauge field. It is interesting to investigate the coupled dynamics of the $B-L$ and $\text{U}(1)_{Y}$ gauge fields and the baryon asymmetry of the universe in this case, although we leave this for future work.

### v2

@@ IV Model with Dirac neutrinos
As we have discussed in the previous section, we need two requirements to generate the baryon asymmetry of the universe. The first requirement is the washout of the BLB-L charges in the standard model sector and the chiral sector. The second requirement is the decoupling of the interactions before the decay of the magnetic helicity. As the Yukawa interaction is decoupled, the first requirement is not satisfied, because the Lagrangian respects the global BLB-L symmetry. However, the Yukawa interaction is not the only possible interaction between the right-handed neutrinos and the standard model leptons. For example, we can introduce higher-dimensional operators between the right-handed neutrinos and the standard model leptons. An example of the relevant operators consistent with the symmetries is a dimension-8 operator,
$\mathcal{L}_{\text{dim-8}}=\frac{1}{\Lambda^{4}}\quantity(HN_{i}^{T}\mathcal{C}L_{j})^{2}+\text{H.c.},$ (IV.68)
@@
where $\Lambda$ is the cutoff scale. In general, the cutoff scale is expected to be the Planck scale, $\Lambda\sim M_{p}$. If we introduce additional heavy fields, the cutoff scale can be much smaller than the Planck scale. Another possibility, as seen in some models called the Dirac seesaw models (e.g., Ref. Roncadelli and Wyler (1983); Ma and Srivastava (2015)), involves the introduction of new heavy particles that equilibrate the BLB-L charges in the standard model sector and chiral sector. As these models introduce additional charged particles and the dynamics of the BLB-L gauge field change, for simplicity, we focus on the simplest model with higher-dimensional operators. With higher-dimensional operators, the BLB-L charges in the standard model sector and the chiral sector are at least partially equilibrated. Let $c$ denote the fraction of the BLB-L charge washed out; $c=1$ if the BLB-L charges are totally equilibrated.
@@
Figure 4: The excluded region of the parameter space in terms of the BLB-L asymmetry and the Hubble constant at the end of the inflation. The left panel is for $g_{B-L}=0.1$ and the right panel is for $g_{B-L}=0.01$. The blue region is the excluded region by the energy density of the magnetic field and the generated Fermions. The green region is the excluded region by the requirement of the magnetic Reynolds number, $R_{m}<1$. We also indicate the contour with $R_{e}=1$ by the green dotted line. The region left to the green dotted line is the region where the kinetic Reynolds number is smaller than unity. The purple, brown, light blue, and yellow dashed lines are the contours of $\xi=5,10,15,20$, respectively.
@@
In Fig. 4, we show the excluded region of the parameter space in terms of the BLB-L asymmetry and the Hubble constant at the end of the inflation, assuming that the magnetic helicity is generated during the axion inflation as discussed in Sec. II. The blue region is the excluded region where the energy density of the magnetic field and the generated asymmetric fermion are larger than $1\,\%$ of the total energy density of the universe at the end of the inflation. The green region is the excluded region by the requirement of the magnetic Reynolds number, $R_{m}\lesssim 1$. We also show the contour with $R_{e}=1$ by the green dotted line as a reference. The region left to the green dotted line is the region where the kinetic Reynolds number is smaller than unity.
@@
From the figure, we can see that we need a larger BLB-L helicity for the magnetic field not to dissipate for a conservative estimate of the magnetic Reynolds number adopted in Sec. II. This requires that the fraction of the BLB-L charge washed out, $c$ is much smaller than unity, as $c\simeq\frac{79}{28}\eta_{0}/\eta_{\cal H}$ to explain the baryon asymmetry of the universe, assuming that the anomalous charge cancellation does not occur. Alternatively, we can dilute the baryon asymmetry by the entropy production after the decay of the magnetic helicity. Another possibility is to consider the case where the helicity is decaying by the dissipation or the anomalous charge cancellation, but the final BLB-L charge is determined by the balance between the decay of the helicity and the washout process. Once more, since the condition adopted here for the magnetic field not to dissipate is a rough estimate, the possibility that the magnetic field survives at smaller $\eta_{H}$ is not completely ruled out Domcke et al. (2023). Determining the parameter space for the successful baryogenesis scenario requires more detailed analyses of the helicity evolution and decay by MHD as well as fermion production, which are beyond the scope of this paper. To conclude this section, we note that the generation of the baryon asymmetry of the universe in the model with Dirac neutrinos can be seen as an implementation of Dirac leptogenesis Dick et al. (2000); Murayama and Pierce (2002). In Dirac leptogenesis, the decay of a heavy particle produces a BLB-L asymmetry in both the standard model sector and the right-handed neutrino sector. In our scenario, this asymmetry is generated through the decay of the magnetic helicity of the BLB-L gauge field.
@@ V Discussion and Conclusion
In this paper, we have proposed a new scenario to generate the baryon asymmetry of the universe. We have discussed the decay of the magnetic helicity of the BLB-L gauge field, which is generated, e.g., during axion inflation, after the BLB-L gauge symmetry is Higgsed. The net BLB-L asymmetry in the standard model sector is the combination of the asymmetry generated during inflation, which is subject to the washout effect, and the asymmetry generated by the BLB-L helicity decay. For both the seesaw model with Majorana neutrinos and the model with Dirac neutrinos, we have investigated if the baryon asymmetry of the universe can be generated in the model through our mechanism. We have shown that the observed baryon asymmetry can be generated in the model with Dirac neutrinos if the effect of the washout process is decoupled before the decay of the magnetic helicity. For the model with Majorana neutrinos, we have shown that the baryon asymmetry of the universe cannot be generated in the minimal seesaw model with the second-order phase transition, although the baryon asymmetry can be generated in more complicated models.
While our discussion on the helicity evolution is based on order-of-magnitude estimates from MHD, the fundamental principles of our scenario remain robust. These principles include the conservation of BLB-L magnetic helicity, the equilibration and subsequent decoupling of BLB-L charges between the standard model sector and the chiral sector, and the decay of magnetic helicity following the BLB-L phase transition. However, to precisely quantify the baryon asymmetry of the universe, detailed analytical formulation and numerical simulations of helicity decay, particularly near the phase transition, are essential.
@@
Finally, in this paper, we have ignored the effect of the $\text{U}(1)_{Y}$ gauge field and and its dynamics. For example, the kinetic mixing between the BLB-L gauge field and the $\text{U}(1)_{Y}$ can generate the helicity of the $\text{U}(1)_{Y}$ gauge field Kamada et al. (2018). Also, if the generated BLB-L charge is large and the temperature is low enough, the chiral plasma instability may generate the helicity of the $\text{U}(1)_{Y}$ gauge field. It is interesting to investigate the coupled dynamics of the BLB-L and $\text{U}(1)_{Y}$ gauge fields and the baryon asymmetry of the universe in this case, although we leave this for future work.